### v1

@@ Abstract
In general-sum stochastic games, a stationary Stackelberg equilibrium (SSE) does not always exist, in which the leader maximizes leader’s return for all the initial states when the follower takes the best response against the leader’s policy. Existing methods of determining the SSEs require strong assumptions to guarantee the convergence and the coincidence of the limit with the SSE. Moreover, our analysis suggests that the performance at the fixed points of these methods is not reasonable when they are not SSEs. Herein, we introduced the concept of Pareto-optimality as a reasonable alternative to SSEs. We derive the policy improvement theorem for stochastic games with the best-response follower and propose an iterative algorithm to determine the Pareto-optimal policies based on it. Monotone improvement and convergence of the proposed approach are proved, and its convergence to SSEs is proved in a special case.
@@ 1 Introduction
A Markov decision process (MDP) is a mathematical framework modeling the decision-making of the agent in dynamic environments, which serves as the theoretical foundation of reinforcement learning (RL). The evaluation value of the agent’s policy is determined by assessing the state-action sequences generated by the policy in the environment with the reward function. This evaluation represents the expectation of cumulative discounted reward. An optimal policy is defined as one that maximizes the evaluation value.
@@
One of the extensions of MDPs for the multi-agent is a stochastic game, also known as a Markov game, where multiple agents with individual reward functions simultaneously try to develop policies that maximize their expected cumulative discounted rewards in one environment. Stochastic games are initially proposed in game theory, and the solution is defined as a tuple of policies that form an equilibrium, such as a Nash equilibrium.The problem setting is characterized by the relationship in the reward functions, such as the zero-sum setting and the fully cooperative setting.
@@
In this study, we addressed the problem of determining a Stackelberg equilibrium in 2-player stochastic games with general relationships in the rewards. A Stackelberg equilibrium is defined as a tuple of strategies of two types of agents, “leader” and “follower,” such that the leader’s strategy maximizes the leader’s payoff when the follower always takes the strategy that is the best response with respect to the follower’s payoff against the leader’s strategy. We aim to find the stationary policy that forms the Stackelberg equilibrium in general-sum stochastic games, that is called SSE. We assume that the follower’s best responses are always computable and the environment is known.
Under an identical setting as ours, Bucarey et al. (2022) proposed an algorithm for finding the SSE. However, this algorithm requires strong assumptions to guarantee convergence to the SSEs. Moreover, this algorithm can converge to low-quality policies when it does not converge to the SSEs Bucarey et al. (2019). As an SSE does not always exist in the general-sum setting Bucarey et al. (2022), the algorithm should converge to the SSE if it exists; otherwise, the algorithm must converge to policies that achieve reasonably good performance. We have developed such an algorithm.
Our contributions are summarized as follows.
First, theoretical analyses of the method proposed by Bucarey et al. (2022) reveal that its fixed point is not necessarily a reasonable policy when it is not guaranteed to converge to the Stackelberg equilibrium. (Section 5.1)
Second, a new sufficient condition for the existence of the SSE includes the one derived in Bucarey et al. (2022). A necessary and sufficient condition for the case of the myopic leader is derived. (Section 5.2)
Third, a concept of the Pareto-optimality is introduced as a reasonable alternative to the SSE. There always exist policies that realize the Pareto-optimal value functions or their neighborhood with arbitrary precision. Moreover, it agrees with the SSE if the equilibrium exists. (Section 6.1)
Lastly, a policy improvement theorem for general-sum SSGs is derived (Section 6.2), based on which a novel algorithm is proposed. Monotone improvement of the state values and convergence of our algorithm are proven in this study (Section 6.3).
@@ 2.1 Markov Decision Process
An MDP is a stochastic process where the state transition depends on the agent’s action and the rewards are defined. MDPs are defined as a tuple $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$: a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, an initial state distribution $\rho\in\Delta(\mathcal{S})$, a bounded reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a discount rate $\gamma\in[0,1)$, where $\Delta(\mathcal{S})$ is a set of probability distributions on $\mathcal{S}$.
The agent determines its action following its Markov policy. Let $\mathcal{W}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ be a set of (stochastic) decision rules, then at the time step $t\in\mathbb{N}$, the agent has the decision rule $f_{t}\in\mathcal{W}$ and selects the action $a_{t}\in\mathcal{A}$ in the current state $s_{t}\in\mathcal{S}$ as $a_{t}\sim f_{t}(s_{t})$. The conditional probability is denoted as $f_{t}(a_{t}|s_{t})$. If the decision rule is equal at any time step ($f_{t}=f$), the Markov policy is called a stationary policy and denoted by a single decision rule $f\in\mathcal{W}$.
@@
for all $s\in\mathcal{S}$, where $\mathbb{E}^{f}\left[\cdot\right]$ is the expectation over $\{(s_{t},a_{t})\}_{t=0}^{\infty}$ generated under $(f,p,\rho)$. $V^{f}(s)$ is called a state value function of $f$.
There always exists a deterministic optimal policy Sutton & Barto (2018) such that, for all $s\in\mathcal{S}$,
$f^{*}(s)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\bigg{\{}r(s,a)\\ +\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\max_{f\in\mathcal{W}}V^{f}(s^{\prime})\right]\bigg{\}}.$ (1)
The main algorithms used to determine the optimal policy are value iteration and policy iteration when the transition function and the reward function are known and RL when they are unknown Sutton & Barto (2018).
@@ 2.2 Stochastic Game and Stackelberg Game
Stochastic games Shapley (1953); Solan (2022), also called Markov games, are the extension of MDPs for the multi-agent. When there are two agents, A and B, they have finite action spaces $\mathcal{A}$ and $\mathcal{B}$, bounded reward functions $r_{A}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$ and $r_{B}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$, discount rates $\gamma_{A}\in[0,1)$ and $\gamma_{B}\in[0,1)$ and sets of decision rules $\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ and $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\}$, respectively. The state $s\in\mathcal{S}$ is shared among the agents, and the transition function is defined as $p:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta(\mathcal{S})$, that is, the transition probability depends on the actions of all the agents.
The expectation of the cumulative discounted reward under stationary policies $f\in\mathcal{W}_{A}$ and $g\in\mathcal{W}_{B}$ is defined for $r_{A}$ and $r_{B}$, respectively. Then, a pair of policies that are Nash equilibrium or other equilibria can be defined as the solutions to the problem by considering each cumulative reward as the payoff of the agent.
The relationship of the reward functions between agents characterizes the game. In game theory, the game where $r_{A}=-r_{B}$ is called zero-sum game, and the others are called non-zero-sum game. Both are called general-sum games. In the context of a multi-agent RL, situations where $r_{A}\propto-r_{B}$ and $r_{A}\propto r_{B}$ are called competitive and cooperative, respectively.
The Stackelberg game is a game model where a leader first selects its strategy and then a follower selects its strategy such that it maximizes the follower’s payoff against the leader’s strategy. This follower’s strategy is called the best response strategy against the leader’s strategy. Then, when the leader’s strategy maximizes the leader’s payoff under the follower’s best response, the pair of strategies is called a Stackelberg equilibrium. In this study, we consider the extension of Stackelberg games to stochastic games, such as stochastic Stackelberg game (SSG). An example of the application of SSG is provided in Appendix M. Unlike the single-agent MDP, the existence of the optimal policy, the SSE policy defined below, is not guaranteed, and the SSE policy is not deterministic even if it exists.
@@ 3 Problem Setting
A general-sum SSG is formulated. Its solutions are defined as the policies that are Stackelberg equilibria.
In a 2-player stochastic game represented by $(\mathcal{S},(\mathcal{A},\mathcal{B}),\hat{p},\rho,(\hat{r}_{A},\hat{r}_{B}),(\gamma_{A},\gamma_{B}))$, agent A is called a “leader” and agent B is called a “follower.” The sets of decision rules are defined by
$\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\},$ $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\},$
and the sets of deterministic decision rules are defined by
$\mathcal{W}_{A}^{d}:=\{f:\mathcal{S}\to\mathcal{A}\},$ $\mathcal{W}_{B}^{d}:=\{g:\mathcal{S}\to\mathcal{B}\}.$
@@
where $\mathbb{E}^{fg}\left[\cdot\right]$ is the expectation over the state-action sequence $\{(s_{t},a_{t},b_{t})\}_{t=0}^{\infty}$ generated under $(f,g,\hat{p},\rho)$.
@@
Let $s\in\mathcal{S}$ be a state and $f\in\mathcal{W}_{A}$ a leader’s decision rule. The set of the follower’s best responses $R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$ against $f$ is defined as
@@
where
$R_{B}^{*}(s,f):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f(s),b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),b)}\left[\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s^{\prime})\right].$
As the optimal policies in single-agent MDPs are represented as in Equation 1, we can observe that the set of best responses $R_{B}^{*}(f)$ is the set of the optimal policies with respect to single-agent MDPs when the leader’s policy is fixed to the stationary policy $f$. It is also shown that $R_{B}^{*}(f)$ is non-empty for all $f\in\mathcal{W}_{A}$.
The follower that always adopts the best response against the leader’s policy is called the “best-response follower.” If the leader adopts a stationary policy $f\in\mathcal{W}_{A}$, the best-response follower adopts the stationary policy $g\in R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$.
We assume that $R_{B}^{*}(s,f)$ is a singleton for all $s\in\mathcal{S},f\in\mathcal{W}_{A}$. It follows that $R_{B}^{*}(f)$ is also singleton for all $f\in\mathcal{W}_{A}$; thus, the follower’s best response against $f$ is unique. For simplicity, let $R_{B}^{*}(f)$ be the best response. This enables us to define a state value function $V_{A}^{fR_{B}^{*}(f)}$ of $f$ under the best-response follower. For simplicity, we denote
@@
Notation
[…]
4 Related Works

### v2

@@ Abstract
We address two-player general-sum stochastic Stackelberg games (SSGs), where the leader’s policy is optimized considering the best-response follower whose policy is optimal for its reward under the leader. Existing policy gradient and value iteration approaches for SSGs do not guarantee monotone improvement in the leader’s policy under the best-response follower. Consequently, their performance is not guaranteed when their limits are not stationary Stackelberg equilibria (SSEs), which do not necessarily exist. In this paper, we derive a policy improvement theorem for SSGs under the best-response follower and propose a novel policy iteration algorithm that guarantees monotone improvement in the leader’s performance. Additionally, we introduce Pareto-optimality as an extended optimality of the SSE and prove that our method converges to the Pareto front when the leader is myopic.
keywords:
Stochastic Stackelberg Games; Policy Iteration; Convergence Analysis.
@@ 1 Introduction
A Markov decision process (MDP) is a mathematical framework that models the decision-making of agents in dynamic environments, which serves as the foundation of reinforcement learning (RL). The performance of the agent’s policy is evaluated by assessing the state-action sequences it produces with the reward function. This performance is quantified by the expected cumulative discounted reward. An optimal policy is one that maximizes this expected return.
@@
One extension of MDPs for multi-agent systems is a stochastic game, also known as a Markov game, where multiple agents with individual reward functions simultaneously attempt to develop policies that maximize their expected cumulative discounted rewards within a single environment. Stochastic games were initially proposed in game theory, and the solution is defined as a tuple of policies that form a specific type of equilibrium, such as a Nash equilibrium. The problem setting is also characterized by the relationship between the reward functions, such as the zero-sum (i.e., competitive) and fully cooperative settings.
@@
In this study, we addressed the problem of determining stationary Stackelberg equilibria (SSEs) in 2-player stochastic games with general relationships in the rewards. An SSE is defined as a tuple of stationary policies of two types of agents, leader and follower, such that the leader’s policy maximizes the leader’s payoff when the follower takes the policy that is the best response to the leader’s policy with respect to the follower’s payoff. We assume that the follower always takes the best response to any leader’s policy. We refer to such a follower as a best-response follower. Additionally, we assume that the follower’s best responses are always computable to the leader and that the environment is known. No other assumptions are made on the follower’s algorithm.
This situation arises, for example, in the design of e-commerce platforms aimed at the maximization of the site owner’s profit. The site user (follower) repeatedly makes purchasing decisions based on their preferences. In contrast, the site owner (leader) can configure various elements of the platform, such as page transitions, advertisements, and an incentive bonus. Leveraging knowledge of the user’s past behavior, the owner seeks to maximize long-term profit by anticipating the user’s responses. Assuming the user’s responses are always optimal for their preferences and the owner’s prediction is accurate enough, the resulting owner’s problem can then be formulated as an SSG with a known best-response follower.
As an SSE does not always exist in the general-sum setting (Bucarey et al., 2022), the algorithm should converge a leader’s policy to the SSE if it exists; otherwise, it must converge to a policy that achieves reasonably satisfactory performance. However, existing methods based on dynamic programming operators (Bucarey et al., 2022; Zhang et al., 2020a) and policy gradient methods with total derivatives (Zheng et al., 2022; Vu et al., 2022) do not guarantee the convergence to the SSEs in our setting, even if they exist. Moreover, these algorithms may converge leaders’ policies to low-quality ones because they do not guarantee monotone improvement of the leaders’ performance under the best-response follower. In this work, we focus on policy improvement methods and develop an algorithm that satisfies the above requirements.
1.1 Main Contributions
Our contributions are summarized as follows. First, we explicitly derive the fixed point of the dynamic programming operator used in Bucarey et al. (2022); Zhang et al. (2020a). This result reveals that the fixed point does not necessarily bring a reasonable leader policy when it is not guaranteed to be an SSE (Section˜5). Second, a policy improvement theorem for general-sum stochastic Stackelberg games with the best-response follower is derived (Section˜6.2), based on which a novel policy iteration algorithm is proposed (Section˜6.3). Finally, we introduce the concept of Pareto-optimality as an extended optimality of the SSE. Policies that realize the Pareto-optimal value functions or their neighborhood with arbitrary precision always exist. Moreover, Pareto-optimality agrees with the SSE it exists (Section˜6.1). Then, we prove that the proposed method monotonically improves the state values toward the Pareto front and converges to the front when the leader is myopic (Section˜6.3). To the best of our knowledge, this is the first theoretical guarantee in general-sum myopic-leader SSGs.
@@ 2.1 Markov Decision Process
An MDP is a stochastic process with rewards in which the state transitions depend on the actions of an agent. An MDP is defined as a tuple $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$: a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, an initial state distribution $\rho\in\Delta(\mathcal{S})$, a bounded reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a discount rate $\gamma\in[0,1)$, where $\Delta(\mathcal{S})$ is a set of probability distributions on $\mathcal{S}$.
The agent determines its action following its Markov policy. Let $\mathcal{W}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ be a set of (stochastic) decision rules. Then, at the time step $t\in\mathbb{N}$, the agent has the decision rule $f_{t}\in\mathcal{W}$ and selects the action $a_{t}\in\mathcal{A}$ in the current state $s_{t}\in\mathcal{S}$ as $a_{t}\sim f_{t}(s_{t})$. The conditional probability is denoted as $f_{t}(a_{t}|s_{t})$. If the decision rule is invariant over time ($f_{t}=f$), the Markov policy is called a stationary policy and is simply denoted by $f\in\mathcal{W}$.
@@
for all $s\in\mathcal{S}$, where $\mathbb{E}^{f}\left[\cdot|s_{0}=s\right]$ is the expectation over $\{(s_{t},a_{t})\}_{t=0}^{\infty}$ generated under $(f,p)$ given initial state $s$. $V^{f}(s)$ is called a state value function of stationary policy $f$. Hereafter, we consider only stationary policies and refer to them simply as policies.
There always exists a deterministic optimal policy (Sutton and Barto, 2018) such that
$f^{*}(s)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\bigg\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\max_{f\in\mathcal{W}}V^{f}(s^{\prime})\right]\bigg\}$ (1)
for all $s\in\mathcal{S}$. The main algorithms used to determine the optimal policy are value iteration and policy iteration when the transition and reward functions are known, and RL when they are unknown (Sutton and Barto, 2018).
@@ 2.2 Stochastic Game and Stackelberg Game
Stochastic games (Shapley, 1953; Solan, 2022), also known as Markov games, are an extension of MDPs for multi-agent settings. When there are two agents, A and B, they have finite action spaces $\mathcal{A}$ and $\mathcal{B}$, bounded reward functions $r_{A}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$ and $r_{B}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$, discount rates $\gamma_{A}\in[0,1)$ and $\gamma_{B}\in[0,1)$ and sets of policies $\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ and $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\}$, respectively. The state $s\in\mathcal{S}$ is shared among the agents, and the transition function is defined as $p:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta(\mathcal{S})$; that is, the transition probability depends on the actions of all the agents.
The expectation of the cumulative discounted reward under stationary policies $f\in\mathcal{W}_{A}$ and $g\in\mathcal{W}_{B}$ is defined for $r_{A}$ and $r_{B}$, respectively. Then, equilibrium policies, such as Nash equilibrium, can be identified as solutions to the problem by considering each cumulative reward as the agents’ payoff.
The relationship of the reward functions between agents characterizes the game. In game theory, a game where $r_{A}=-r_{B}$ is called a zero-sum game, and the others are called non-zero-sum games. Both are called general-sum games. In the context of a multi-agent RL, situations where $r_{A}\propto-r_{B}$ and $r_{A}\propto r_{B}$ are called competitive and cooperative, respectively.
The Stackelberg game is a game model in which a leader selects its strategy first, and then a follower selects its strategy that maximizes the follower’s payoff against the leader’s strategy. This follower’s strategy is called the best response strategy against the leader’s strategy. When the leader’s strategy maximizes the leader’s payoff under the follower’s best response, the pair of strategies is called a Stackelberg equilibrium. In this study, we consider the extension of Stackelberg games to stochastic games, such as the stochastic Stackelberg game (SSG). Unlike the single-agent MDP, the optimal policy, SSE policy defined below, is not guaranteed to exist, and even if it does, it is not necessarily deterministic. An example scenario where the SSE policy does not exist is provided in the next section.
@@ 3 Problem Setting
A general-sum SSG is represented by $(\mathcal{S},(\mathcal{A},\mathcal{B}),\hat{p},\rho,(\hat{r}_{A},\hat{r}_{B}),(\gamma_{A},\gamma_{B}))$, where the subscripts $A$ and $B$ indicate the leader and follower, respectively. The sets of policies are defined by
$\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\},\quad\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\},$
and the sets of deterministic policies are defined by
$\mathcal{W}_{A}^{d}:=\{f:\mathcal{S}\to\mathcal{A}\},\quad\mathcal{W}_{B}^{d}:=\{g:\mathcal{S}\to\mathcal{B}\}.$
@@
[…]
@@
Let $s\in\mathcal{S}$ be a state and $f\in\mathcal{W}_{A}$ a leader’s policy. The set of the follower’s best responses $R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$ against the leader’s policy $f$ is defined as
@@
where the best response action at state $s$ is defined as
$R_{B}^{*}(s,f):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f(s),b)+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),b)}\left[\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s^{\prime})\right].$
As the optimal policies in single-agent MDPs are represented as in Equation˜1, we can observe that the set of best responses $R_{B}^{*}(f)$ with fixed leader’s stationary policy $f$ is the set of the optimal policies in terms of single-agent MDPs. It is also shown that $R_{B}^{*}(f)$ is non-empty for all $f\in\mathcal{W}_{A}$.
We assume that the follower is a best-response follower, meaning that it always adopts the best response to the leader’s policy. If the leader adopts a stationary policy $f\in\mathcal{W}_{A}$, the best-response follower adopts the stationary policy $g\in R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$. This setup is essentially equivalent to a situation where we can only control the leader’s policy, and the follower independently learns its policy with the ability to find the best responses in a finite amount of time for any leader’s policy. No other assumptions are made about the follower’s learning process.
We assume that the set of the follower’s best responses, $R_{B}^{*}(s,f)$, is a singleton for each $s\in\mathcal{S}$ and for each $f\in\mathcal{W}_{A}$. Otherwise, we break ties deterministically. It follows that $R_{B}^{*}(f)$ is also a singleton for all $f\in\mathcal{W}_{A}$; thus, the follower’s best response against $f$ is unique. For simplicity, let $R_{B}^{*}(f)$ be the best response. This enables us to define a state value function $V_{A}^{fR_{B}^{*}(f)}$ of $f$ under the best-response follower. For simplicity, we denote
@@
Game without SSEs
We provide an example game where SSEs do not exist. This example is first introduced by Jean-Marie in his seminar talk (Jean-Marie et al., 2022) related to Bucarey et al. (2022). We consider a game with two states $\mathcal{S}=\{s_{1},s_{2}\}$, two leader’s actions $\mathcal{A}=\{a_{1},a_{2}\}$, and two follower’s actions $\mathcal{B}=\{b_{1},b_{2}\}$. The dynamics and the reward signals are deterministic and represented in Figure˜1. Let the leader’s policy be denoted as $f_{s_{1}}(a_{1})=p\in[0,1]$ and $f_{s_{2}}(a_{1})=q\in[0,1]$.
Given the leader’s policy, it reduces to a standard MDP for the follower. The optimal value function for the follower is the solution to
$V_{B}^{*}(s_{1})$ $=\max\{\gamma_{B}V_{B}^{*}(s_{1}),x+\gamma_{B}V_{B}^{*}(s_{2})\}\cdot p+(-y+\gamma_{B}V_{B}^{*}(s_{2}))\cdot(1-p);$
$V_{B}^{*}(s_{2})$ $=\max\{\gamma_{B}V_{B}^{*}(s_{2}),x+\gamma_{B}V_{B}^{*}(s_{1})\}\cdot q+(-y+\gamma_{B}V_{B}^{*}(s_{1}))\cdot(1-q).$
There are three cases depending on the leader’s policy $f$, summarized in Table˜1. If $\gamma_{B}y>x$, we know that $(p,q)=(1,0)$ leads to the case of $\gamma_{B}V_{B}^{*}(s_{1})>x+\gamma_{B}V_{B}^{*}(s_{2})$. Then, the maximum value of the leader’s value at state $s_{1}$ is obtained as $V_{A}^{f{\dagger}}(s_{1})=\frac{1}{1-\gamma_{A}}$. Similarly, we can check that $(p,q)=(0,1)$ leads to the case of $\gamma_{B}V_{B}^{*}(s_{2})>x+\gamma_{B}V_{B}^{*}(s_{1})$. Then, the maximum value of the leader’s value at state $s_{2}$ is obtained as $V_{A}^{f{\dagger}}(s_{2})=\frac{1}{1-\gamma_{A}}$. We can see that there is a tradeoff between the values at $s_{1}$ and $s_{2}$ and two Pareto optimal points exist: $(V_{A}^{f{\dagger}}(s_{1}),V_{A}^{f{\dagger}}(s_{2}))=(\frac{1}{1-\gamma_{A}},\frac{\gamma_{A}}{1-\gamma_{A}})$ and $(\frac{\gamma_{A}}{1-\gamma_{A}},\frac{1}{1-\gamma_{A}})$. That is, there does not exist a single optimal policy that maximizes the values at all states simultaneously, implying that an SSE does not exist.
Table 1: The follower’s best response and the leader’s value under the best response follower.
Case $R_{B}^{*}(s_{1},f),R_{B}^{*}(s_{2},f)$ $V_{A}^{f{\dagger}}(s_{1})$, $V_{A}^{f{\dagger}}(s_{2})$
$\gamma_{B}V_{B}^{*}(s_{1})>x+\gamma_{B}V_{B}^{*}(s_{2})$ $b_{1},b_{2}$ $\frac{p}{1-\gamma_{A}},\frac{p\gamma_{A}}{1-\gamma_{A}}$
$\gamma_{B}V_{B}^{*}(s_{2})>x+\gamma_{B}V_{B}^{*}(s_{1})$ $b_{2},b_{1}$ $\frac{q\gamma_{A}}{1-\gamma_{A}},\frac{q}{1-\gamma_{A}}$
$\lvert V_{B}^{*}(s_{1})-V_{B}^{*}(s_{2})\rvert<x/\gamma_{B}$ $b_{2},b_{2}$ $0,0$
$s_{1}$$s_{2}$($a_{1}$, $b_{1}$) / (1, 0)($a_{1}$, $b_{1}$) / (1, 0)($a_{1}$, $b_{2}$) / (0, $x$) ($a_{2}$, *) / (0, y-y)
Figure 1: An example of a game without SSEs. Transitions are deterministic and represented by arrows. Each label shows (leader’s action, follower’s action) /  (leader’s reward, follower’s reward), where $x>0$ and $y>0$.
4 Related Work